OpenStudy (anonymous):

S=4*pi*(r^2) and V=(4*pi*(r^3))/3 find dS/dV

OpenStudy (unklerhaukus):

\[S=4\pi r^2\qquad\frac{\text dS}{\text dr}=8\pi r\] \[V=\frac{ 4 \pi r^3}3\qquad\frac{\text dV}{\text dr}=4\pi r^2\] \[\frac{\text dS}{\text dV}=\frac{\text dS}{\text dr}\cdot\frac{\text dr}{\text dV}=8\pi r\cdot\frac1{4\pi r^2}=\frac 2r\]

OpenStudy (anonymous):

S=4*pi*(r^2) and V=(4*pi*(r^3))/3 a spherical balloon expands as air is pumped into it. what is the rate of increase in the balloon's surface area with volume when its volume is 30(cm^3)

OpenStudy (unklerhaukus):

solve for r, plug r into dS/dV

OpenStudy (anonymous):

can u pls solve it for me